NottheMonga2087:

What is the value of x in the equation below? 6(x-8)=72 12 15 20 24

DumbFounded:

ok so u always multiply by the number in the brackets

NottheMonga2087:

yes multiplication property

NottheMonga2087:

of equality

DumbFounded:

so use the answer choices and answer the question

DumbFounded:

6(12-8)=72 6(15-8)=72 6(20-8)=72 6(24-8)=72

DumbFounded:

try thos

DumbFounded:

tell me which one makes more sense

NottheMonga2087:

to find the answer would you use what times what equals 72

DumbFounded:

first you would do this 6x-48=72

DumbFounded:

then you would add 48 to both sides of the equation

NottheMonga2087:

ok

DumbFounded:

then what would you get?

NottheMonga2087:

120

DumbFounded:

good! now then you would do 6x/6 = 120/6

DumbFounded:

and you get?

NottheMonga2087:

20!

DumbFounded:

good job!

DumbFounded:

thats the correct answer

NottheMonga2087:

thank you very much

DumbFounded:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @NottheMonga2087 thank you very much \(\color{#0cbb34}{\text{End of Quote}}\) mp chief

DumbFounded:

*np